$SPAD/src/input telford.input

T.Doug Telford December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool telford.output
)set message test on
)set message auto off
)clear all
--S 1 of 5
p(0)==1
--R                                                                   Type: Void
--E 1
--S 2 of 5
p(1)==x
--R                                                                   Type: Void
--E 2
--S 3 of 5
p(n)==((2*n-1)*x*p(n-1)-(n-1)*p(n-2))/n
--R                                                                   Type: Void
--E 3
--S 4 of 5
p(10)
--R   Compiling function p with type Integer -> Polynomial(Fraction(
--R      Integer))
--R   Compiling function p as a recurrence relation.
--R
--R        46189  10   109395  8   45045  6   15015  4   3465  2    63
--R   (4)  ----- x   - ------ x  + ----- x  - ----- x  + ---- x  - ---
--R         256          256        128        128        256      256
--R                                          Type: Polynomial(Fraction(Integer))
--E 4
--S 5 of 5
coefficient(p(10),x,10)
--R
--R        46189
--R   (5)  -----
--R         256
--R                                          Type: Polynomial(Fraction(Integer))
--E 5
)spool
)lisp (bye)

———-

References^{[1] nothing}

4